Two disks of diameter d are spaced a distance r ap
Answer #1

(a) Because the electric field is constant, V = Electric field*distance = E*r

(b) The capacitance of the capacitor is C = ε0*area/distance = ε0*π(d^2/4)/r =  ε0*π*d^2/4r

The magnitude of the charge on each disk is therefore

Q = C∆V = [ ε0*π*d^2/4r]*E*r = ε0*E*π*d^2/r